# Simulations of gravitational collapse in null coordinates IV: evolving through the event horizon, with an application to the spherical charged scalar field

### v1

@@ I Introduction
I.1 The problem posed by evolving through an event horizon in null coordinates
In this series of papers, starting with axinull_formulation , we have been developing numerical simulations of gravitational collapse on null cones emanating from a regular centre: in the present paper, we examine formulations that allow us to evolve from regular data through the black hole event horizon. After this Introduction, we will restrict to spherical symmetry in 3+1 spacetime dimensions, but the ideas generalise straightforwardly, and will be applied to non-spherical critical collapse elsewhere.
In Paper II of the series axinull_critscalar , we investigated the threshold of black hole formation for a massless scalar field coupled to general relativity in axisymmetry, adapting the coordinates to the expected self-similar critical solution. While, at moderate non-sphericity, this allowed us to fine-tune to the threshold of black hole formation to machine precision, without any need for adaptive mesh-refinement, we could not reach non-sphericities as high as others had reached before us. In hindsight, this was due both to our choice of radial gauge (how the radial coordinate $x$ labels points on each null cone generator) and our choice of formulation (which of the Einstein equations are used in the code and which are not). For our notation for the metric, see already Eq. (1) below.
@@ I.1 The problem posed by evolving through an event horizon in null coordinates
This paper presents a second formulation, where the Raychaudhuri equation is solved for $R$, rather than for $G$. In the literature, this approach has been used only together with the affine radial gauge $G=1$. Here we evolve $G$, which allows us to use an arbitrary radial shift $B$, and in particular one adapted to critical collapse. We also present a third formulation where both $R$ and $G$ are evolved, and the Raychaudhuri equation is not used. In the literature, this has been used only with double null gauge $B=0$, but $B$ can in fact be chosen freely. Here we point out that all three formulations can be combined with any radial gauge choice. We present extended tests of these three formulations in spherical symmetry, with charged scalar field matter.
@@
I.2 Formulations and radial gauges in null coordinates
We now give a more precise statement of what we just said about formulations, but without restricting the spacetime dimension or symmetry. (In the remainder of the paper, we will restrict to spherical symmetry in 3+1 dimensions.) We remind the reader that we consider line elements of the form
@@ I.2 Formulations and radial gauges in null coordinates
(The quantity $B$ replaces the metric coefficient $H:=-g_{uu}$ that we used in Papers I–III. This is just for convenience of notation.) The spacetime dimension is $n+2$, with the $\theta^{i}$ coordinates on an $n$-dimensional manifold, typically $S^{n}$. From this form of $g_{\mu\nu}$ it follows that $g^{uu}=g^{ui}=0$, and $g^{ux}=-G^{-1}$. Hence the surfaces of constant $u$ are null surfaces, and their affinely parameterised generators have tangent vector
@@
We must have $G>0$ for the metric to be regular. $R$ is defined by restricting $\gamma_{ij}$ to have a fixed determinant (such as its value in Minkowski spacetime). The remaining gauge freedom in this ansatz for the line element is to re-label $u$ as $u\to u^{\prime}(u)$, and to relabel the radial coordinate $x$ as $x\to x^{\prime}(u,x,\theta^{i})$. We call a choice of coordinate $x$ a radial gauge.
@@
(See also Paper I axinull_formulation for the full definition of $\rho_{+}$.) The quantity
@@
From (8), with $K/G\geq 0$, we see that once $dR/d\lambda$ has become negative, $R$ reaches zero in finite affine parameter. Equivalently, from (9) we see that once $\rho_{+}$ has become negative, it blows up in finite affine parameter. We can now make the above statement precise: In a non-spherical collapse spacetime, generically there will be $u_{0}<u_{1}$ such that for any coordinate null cone with $u_{0}<u<u_{1}$, the expansion $\rho_{+}$ remains positive on some of its generators, while on others it becomes negative and then blows up. This could be because it has reached a curvature singularity, or because the null cone has developed a caustic.
We now come to the role of the Raychaudhuri equation in numerical relativity on null cones. In its form (6) and with $G$ given, it can be solved as a second-order ordinary differential equation (ODE) in $x$ for $R$. In its form (7) and with $R$ given, it can be solved as a first-order ODE for $G$. We also have wave equation-like evolution equations $G_{,ux}=$... and $R_{,ux}=$... for both $G$ and $R$. This gives rise to three formulations, each of which admits an arbitrary radial gauge:
@@
I.3 Plan of this paper
In the following, we will restrict to spherical symmetry in 3+1 dimensions, with charged scalar field matter, but the ideas illustrated here are easily generalised to less symmetry, other spacetime dimensions, and to vacuum or other matter.
We present an implementation of all three formulations set out above in one code, and test them in two settings. In the first of these, we evolve on outgoing null cones that emanate from a regular central worldline. To fix the radial gauge in the initial data, we either set $R=x$ or $G=1$ at $u=0$. We continue in “shifted double-null gauge”, an incremental gauge that is adapted to critical collapse in spherical symmetry (see Sec. III.C of Paper I for a definition and PortoGundlach2022 for a previous application). As a physics application, we compute the fine-structure of the mass and charge scaling laws in critical collapse, which has not been computed before. In the second setting the left boundary $x=0$ is not a regular centre but an ingoing null cone, so that the computational domain is a null rectangle. We continue the initial gauge choice of $R_{x,}=1$ or $G=1$ as double-null gauge $B=0$.
@@ I.3 Plan of this paper
II Field equations
@@ II.1 Field equations in covariant form
The equations admit the gauge freedom $\phi\to e^{iq\alpha}\phi$, $A_{a}\to\nabla_{a}\alpha$ for an arbitrary scalar function $\alpha$, leaving $F_{ab}$ and $D_{a}\phi$, and hence $j^{a}$ and $T_{ab}$, invariant.
@@ II.2 Field equations in spherical symmetry in null coordinates
$\alpha(u,x)=\int_{0}^{u}A_{u}(u^{\prime},0)\,du^{\prime}-\int_{0}^{x}A_{x}(u,x^{\prime})\,dx^{\prime}.$ (21)
@@
As $R\geq 0$ and $G>0$, $\rho_{+}$ has the same sign as $R_{,x}$. Hence a symmetry sphere is trapped if and only if $R_{,x}<0$, and marginally trapped if and only if $R_{,x}=0$. (By trapped, we always mean outer-trapped.)
@@
Hence there is another conserved mass curent analogous to (29).
@@
In initial data on an outgoing null cone with regular centre, we always have $\Xi R<0$ and hence $XR<0$ on our initial data. From (49) we see that $XXR\leq 0$, so $XR<0$ and hence $\Xi R<0$ on the entire domain of dependendence of these initial data. Hence, from (33), a symmetry sphere is trapped if and only if $C>1$ and marginally trapped if and only if $C=1$. This is not obvious from (42), whose right-hand side becomes negative when $R<|Q|$, and so, as we solve (42) but not (49) in algorithm, $\Xi R<0$ may be violated by numerical error.
@@
regardless of spacetime dimension or symmetry.
@@ III.2 Boundary conditions at $x=0$
Where $R=x=0$ is a regular central worldline, $x=0$ is a coordinate singularity (and we omit it from the numerical grid), so we cannot impose simple boundary conditions there. To start up the integration of the hiercharchy equations, we expand in powers of $x$ instead.
At a given $u$, we fit the evolved quantities $\psi$, $\chi$, $R/x$ and/or $G$ either to linear functions of $x$, such as $\psi\simeq\psi_{(0)}+\psi_{(1)}x$, or to quadratic ones, such as $\psi\simeq\psi_{(0)}+\psi_{(1)}x+\psi_{(2)}x^{2}$. (The notation here is compatible with our notation $\psi_{l(n)}(u)$ in Papers I and II, where $l$ denotes the spherical harmonic and $n$ the power of $x$.)
@@
The full expressions are given in Appendix C.
On the other hand, if $x=0$ is a timelike or null inner boundary, we can specify boundary conditions for the hierarchy equations there, subject to constraints. For the case where $x=0$ is an ingoing null cone, we discuss this in detail in Appendix E.
@@ III.3 Time evolution algorithm
We use the “method of lines” framework, where one discretises separately in space and time. We break up each sub-timestep into the following steps, where Steps 1 and 3 depends on the formulation.
Step 0: Assume that on a null cone of constant $u$, the functions $\psi(x)$, $\chi(x)$, and $\ln G(x)$ and/or $R(x)$ are given. When $G$ and $R$ are both given, they must obey the constraint (40). With a regular centre, we fit these functions to power-series expansions in $x$ about $x=0$.
@@
Step 1c: Alternatively, in free evolution, we do nothing. From numerical error, both in the evolution of $R$ and $G$, and in the fitting, the fitting coefficients of $R$ and $G$ will not be exactly compatible with the Raychaudhuri equation, so we need to make an arbitrary choice which set of fitting coefficients we choose for expanding the solution of the hierarchy equations at the centre. The choice made in Appendix C works.
@@
Step 3c: Alternatively, in the other formulations we set $B$ freely. For critical collapse in spherical symmetry, we use sdn gauge (60). On a null rectangle we use double-null gauge $B=0$.
@@
Step 5: Using the method of lines, we evolve $\psi$, $\chi$, and one or both of $R$ and $\ln G$, from $u$ to $u+\Delta u$. This returns us to Step 0.
@@ IV.1 Discretization
This scheme is by construction second-order accurate, and is causal and explicit in the sense that the vector of variables $g_{i}$ depends only on $g_{j}$ with $j<i$ and $f_{j}$ with $j\leq i$: information propagates along a null cone of constant $u$ only towards increasing $x$.
@@
When we solve (64,65) for $(V,R)$, the coefficients $G$ and $K$ are already known.
@@
For comparison, we also use the second or fourth-order Adams-Bashforth method. These need the vector of right-hand sides $F_{i}$ at grid points. We discrete derivatives on grid points as
@@
When we solve (39) for $G$, we first find $\ln V$, where $V:=R_{,x}/G$, on grid points by integration. (In Papers I-III, we used the variables $g:=1/V$ and $\gamma:=\ln g$. We have replaced them with $V$ as they are defined only for $R_{,x}>0$.) We then find $V$ at grid points and hence at mid-points, and hence $G=R_{,x}/V$ at mid-points in a causal and second-order accurate way, given that $R_{,x}$ is known to second order and causally at mid-points.
@@
As $G$ varies rapidly in $u$ and $x$, but is constrained to be strictly positive, we evolve $\ln G$ rather than $G$. We find $G$ at grid points by exponentiating, and hence on mid-points by linear interpolation. To evaluate the compactness $C$ when $G$ is evolved, we us $R_{,x}$ on gridpoints, which is not causal. Again we do not see an adverse effect.
@@ IV.2 Quantities that scale as a power of $x$ at the centre
The quantity $Q$ is obtained by integration of (27). From its definition, the quantity $Q/R^{3}$, representing an average charge density, is expected to be smooth, and to be non-zero at the centre. Numerically, this is of course not guaranteed. We can improve regularity by discretizing
@@ IV.3 Choice of time step
for some $C_{2}\lesssim 1$. In all our gauges $R_{,u}=0$ at $R=0$, and so (77) and (78) are equivalent at the centre.
Since Paper II, we have realised that using $\Delta u=\Delta u_{1}$ may be an unnecessarily small timestep. We have found that $\Delta u=\Delta u_{2}$ with $C_{1}=0.1$ also leads to stable evolutions up to close to the apparent horizon, but with fewer timesteps overall.
@@
If we want to evolve through a horizon, where $R_{,x}=0$, we cannot use the numerical time step $\Delta u_{1}$. In double-null gauge $B=0$, the time step $\Delta u_{2}$ is not defined. The time steps $\Delta u_{3}$ and $\Delta u_{4}$ cannot be used for convergence testing as $\Delta u_{3}$ is independent of $\Delta x$. As an alternative to all these, for example if we want to test convergence of a solution where $R_{,x}=0$ occurs, we can also use the simple timestep

### v2

@@ I Introduction
I.1 Overview
In this series of papers, starting with axinull_formulation , we have been developing numerical simulations of gravitational collapse on null cones emanating from a regular centre. Null coordinates are an established tool in spherical symmetry, but our main motivation is simulating critical collapse on null cones beyond spherical symmetry. As long as no caustics form, null coordinates are ideal for type-II critical collapse because they allow us to adapt the coordinate system to the approximate self-similarity and large range of scales seen there.
In the present paper, we introduce new formulations of the Einstein equations in null coordinates that allow us to evolve from regular data through the black hole event horizon, while keeping the radial coordinate arbitrary. The motivation for this will be given in more detail in the next Subsection, but, in brief, we want to evolve through horizons to accurately characterise black holes, while we also want to keep the radial coordinate arbitrary in order to adapt the coordinates to type-II critical collapse.
To make the presentation more concrete, after this Introduction we apply the new formulations to the spherical Einstein-Maxwell-charged scalar field system. After demonstrating convergence, we turn to type-II critical collapse in this system, and obtain some new physis results. In particular, we verify the critical exponents for the black hole mass and charge, calculate the mass and charge fine structures, and present preliminary evidence that the charge fine structure is not universal. We plan to return to non-spherical type-II critical collapse elsewhere.
The spherical Einstein-Maxwell-charged scalar systems is of current interest also because of the stable extremal critical collapse conjecture of Kehle and Unger KehleUnger22 ; KehleUnger24 . This says that there is an open region of solution space in which the threshold of black hole formation consists of extremal black holes.
In this context, we will need to evolve initial data posed on an outgoing and an ingoing null cone that meet at 2-sphere. The domain of dependence of these data is then a “null rectangle” in the symmetry-reduced spacetime. We present an algorithm for this setup, and demonstrate convergence in selected example solutions. We plan to investigate extremal critical collapse in this system elsewhere.
I.2 The problem posed by evolving through an event horizon in null coordinates
In Paper II of the series axinull_critscalar , we investigated the threshold of black hole formation for a massless scalar field coupled to general relativity in axisymmetry, adapting the coordinates to the expected self-similar critical solution. While, at moderate non-sphericity, this allowed us to fine-tune to the threshold of black hole formation down to machine precision, without any need for adaptive mesh-refinement, we could not reach non-sphericities as high as others had reached before us. In hindsight, this was due both to our choice of radial gauge (how the radial coordinate $x$ labels points on each null cone generator) and our choice of formulation (which of the Einstein equations are used in the code and which are not). For our notation for the metric, see already Eq. (1) below.
@@ I.2 The problem posed by evolving through an event horizon in null coordinates
This paper presents a second formulation, where the Raychaudhuri equation is solved for $R$, rather than for $G$. In the literature, this approach has been used only together with the affine radial gauge $G=1$. Here we evolve $G$, which allows us to use an arbitrary radial shift $B$, and in particular one adapted to critical collapse.
We also present a third formulation, where both $R$ and $G$ are evolved and the Raychaudhuri equation is used only to construct initial data. In the literature, this formulation has been used only with double null gauge $B=0$, but $B$ can in fact be chosen freely.
One of our main points is the observation that all three formulations can be combined with any radial gauge choice.
@@
I.3 Formulations and radial gauges in null coordinates
We now give a more precise statement of what we just said about formulations. For clarity, we do not fix the spacetime dimension or symmetry. (In the remainder of the paper, we will restrict to spherical symmetry in 3+1 dimensions.) We consider line elements of the form
@@ I.3 Formulations and radial gauges in null coordinates
(The quantity $B$ replaces the metric coefficient $H:=-g_{uu}$ that we used in Papers I–III, with $B:=H/(2G)$. This is just for convenience of notation.) The spacetime dimension is $n+2$, and the $\theta^{i}$ are coordinates on an $n$-dimensional manifold, typically $S^{n}$. From this form of $g_{\mu\nu}$ it follows that $g^{uu}=g^{ui}=0$, and $g^{ux}=-G^{-1}$. Hence the surfaces of constant $u$ are null surfaces, and their affinely parameterised generators have tangent vector
@@
We must have $G>0$ for the metric to be regular. $R$ is defined by restricting $\gamma_{ij}$ to have a fixed determinant (such as its value on a unit sphere $S^{n}$ in standard coordinate $\theta^{i}$). The remaining gauge freedom in this ansatz for the line element is to re-label $u$ as $u\to u^{\prime}(u)$, and to relabel the radial coordinate $x$ as $x\to x^{\prime}(u,x,\theta^{i})$. We call a choice of coordinate $x$ a radial gauge.
@@
(See also Paper I axinull_formulation for the definition of $\rho_{+}$ from first principles.) The quantity
@@
From (8), with $K\geq 0$, we see that once $dR/d\lambda$ has become negative, $R$ reaches zero in finite affine parameter. Equivalently, from (9) we see that once $\rho_{+}$ has become negative, it blows up in finite affine parameter. We can now make the above statement precise: In a non-spherical collapse spacetime, generically there will be $u_{0}<u_{1}$ such that for any coordinate null cone with $u_{0}<u<u_{1}$, the expansion $\rho_{+}$ remains positive on some of its generators, while on others it becomes negative and then blows up. This could be because the generator has reached a curvature singularity, or because the null cone has developed a caustic.
We now come to the role of the Raychaudhuri equation in numerical relativity on null cones. In its form (6) and with $G$ given, it can be solved as a second-order ordinary differential equation (from now on, ODE) in $x$ for $R$. In its form (7) and with $R$ given, it can be solved as a first-order ODE for $G$. We also have wave equation-like evolution equations $G_{,ux}=$... and $R_{,ux}=$... for both $G$ and $R$. This gives rise to three formulations, each of which admits an arbitrary radial gauge:
@@
One can retain some of the stability advantages of a maximally constrained formulation in the eR formulation by restricting to the family of “local shifted Bondi” (lsB) radial gauges, where $R=R(u,x)$. We did this in Paper II. Similarly, in the eG formulation, we can restrict to “local shifted affine” (lsa) gauges, where $G=G(u,x)$, to make it almost maximally constrained.
@@
I.4 Plan of this paper
In the remainder of this paper, we will restrict to spherical symmetry in 3+1 dimensions, with charged scalar field matter, but the ideas illustrated here are in principle easily generalised to less symmetry, other spacetime dimensions, and to vacuum or other matter.
We present an implementation of the five formulations described above in one code, and we test the three formulations that admit a general radial gauge in two settings. In the first of these, we evolve on outgoing null cones that emanate from a regular central worldline. To fix the radial gauge in the initial data, we either set $R=x$ or $G=1$ at $u=0$. We continue in “shifted double-null gauge”, see Eq. (60) below, an incremental gauge that is adapted to critical collapse in spherical symmetry (see also Sec. III.C of Paper I for a definition and PortoGundlach2022 for a previous application). As a physics application, we compute the fine-structure of the mass and charge scaling laws in critical collapse, which has not been computed before. In the second setting the left boundary $x=0$ is not a regular centre but an ingoing null cone, so that the computational domain is a null rectangle. We continue the initial gauge choice of $R_{x,}=1$ or $G=1$ as double-null gauge $B=0$.
We do not present any results in the affine and Bondi formulations, as the affine and Bondi gauges that they require do not allow us to make the outer boundary future spacelike (in the first setting) or ingoing null (in the second setting).
@@ I.4 Plan of this paper
II The Einstein-Maxwell-charged scalar system
@@ II.1 Field equations in covariant form
The equations admit the gauge freedom $\phi\to e^{iq\alpha}\phi$, $A_{a}\to A_{a}+\nabla_{a}\alpha$ for an arbitrary scalar function $\alpha$, leaving $F_{ab}$ and $D_{a}\phi$, and hence $j^{a}$ and $T_{ab}$, invariant.
@@ II.2 Field equations in spherical symmetry in null coordinates
$\alpha(u,x)=-\int_{0}^{u}A_{u}(u^{\prime},0)\,du^{\prime}-\int_{0}^{x}A_{x}(u,x^{\prime})\,dx^{\prime}.$ (21)
@@
Hence there is another conserved mass current analogous to (29).
@@
In initial data on an outgoing null cone with regular centre, we always choose $\Xi R<0$ and hence $XR<0$ on $u=0$. From (49) we see that $XXR\leq 0$, so $XR<0$ and hence $\Xi R<0$ on the entire domain of dependendence of these initial data. This is not obvious from (42), whose right-hand side becomes negative when $R<|Q|$, and so, as we solve (42) but not (49), $\Xi R<0$ may be violated by numerical error.
As $R\geq 0$, $G>0$ and $\Xi R<0$, $\rho_{+}$ given above in (3) has the same sign as $R_{,x}$. Hence a symmetry sphere is outer-trapped if and only if $R_{,x}<0$, or equivalently $C>1$.
@@
regardless of spacetime dimension or symmetry, and so $v$ is advected in $x$ with speed B-B.
@@ III.2 Boundary conditions at $x=0$
When $x=0$ is the regular central worldline $R=0$, $x=0$ is a coordinate singularity, and we cannot impose simple boundary conditions there. To start up the integration of the hiercharchy equations, we expand in powers of $x$ instead.
At a given $u$, we fit the evolved quantities $\psi$, $\chi$, $R/x$ and/or $G$ either to linear functions of $x$, such as $\psi\simeq\psi_{(0)}+\psi_{(1)}x$, or to quadratic ones, such as $\psi\simeq\psi_{(0)}+\psi_{(1)}x+\psi_{(2)}x^{2}$. (The notation here is compatible with our notation $\psi_{l(n)}$ in Papers I and II, where $l$ denotes the spherical harmonic and $n$ the power of $x$.)
@@
The full expressions for the expansion coefficients are given in Appendix C.
On the other hand, if $x=0$ is a timelike or null boundary, we can specify boundary conditions for the hierarchy equations there, subject to constraints. For the case where $x=0$ is an ingoing null cone, we discuss this in detail in Appendix E.
@@ III.3 Time evolution algorithm
We use the “method of lines” framework, where one discretises separately in space and time. We break up each sub-timestep into the following steps, where Steps 1 and 3 depend on the formulation. For completeness, we include the affine and Bondi formulations in the following pseudo-code, even though we do not use them for any numerical tests in this paper.
Step 0: Assume that on a null cone of constant $u$, the functions $\psi(x)$, $\chi(x)$, and $G(x)$ and/or $R(x)$ are given. When $G$ and $R$ are both given, they must obey the constraint (40). With a regular centre, we fit these functions to power-series expansions in $x$ about $x=0$.
@@
Step 1c: Alternatively, in the fe formulation we do nothing. From numerical error, both in the evolution of $R$ and $G$, and in the fitting, the fitting coefficients of $R$ and $G$ will not be exactly compatible with the Raychaudhuri equation, so we need to make an arbitrary choice which set of fitting coefficients we choose for expanding the solution of the hierarchy equations at the centre. The choice made in Appendix C works.
@@
Step 3c: Alternatively, in the eR, eG and fe formulations we set $B$ freely. For critical collapse in spherical symmetry, we use sdn gauge (60). On a null rectangle we use double-null gauge $B=0$.
@@
Step 5: Using an ODE integrator at each grid point (the method of lines), we evolve $\psi$, $\chi$, and one or both of $R$ and $\ln G$, from $u$ to $u+\Delta u$. This returns us to Step 0.
@@ IV.1 Discretization
This scheme is explicit, is second-order accurate by construction, and is causal in the sense that the vector of variables $g_{i}$ depends only on $g_{j}$ with $j<i$ and $f_{j}$ with $j\leq i$: information propagates along a null cone of constant $u$ only towards increasing $x$.
@@
When we solve (65,64) for $(V,R)$, the coefficients $G$ and $K$ are already known.
@@
For comparison, we also use the second or fourth-order Adams-Bashforth method. These need the vector of right-hand sides $F_{i}$ at grid points. We discretize $x$-derivatives on grid points as
@@
When we solve (39) for $G$, we first find $\ln V$ on grid points by integration, where $V:=R_{,x}/G$. (In Papers I-III, we used the variables $g:=1/V$ and $\gamma:=\ln g$. We have replaced them with $V$ as they are defined only for $R_{,x}>0$.) We then find $V$ at grid points and hence at mid-points, and hence $G=R_{,x}/V$ at mid-points in a causal and second-order accurate way, given that $R_{,x}$ is known to second order and causally at mid-points.
@@
$G$ varies rapidly in $u$ and $x$, but is constrained to be strictly positive. To avoid it becoming negative through numerical error, we evolve $\ln G$ rather than $G$. We find $G$ at grid points by exponentiating, and hence on mid-points by linear interpolation. To evaluate the compactness $C$ when $G$ is evolved, we us $R_{,x}$ on gridpoints, which is not causal. Again we do not see an adverse effect.
@@ IV.2 Quantities that scale as a power of $x$ at the centre
The quantity $Q$ is obtained by integration of (27). From its definition, the quantity $Q/R^{3}$, representing an average charge density, is expected to be smooth, and to be non-zero at the centre. Numerically, this smoothness is not guaranteed. We can improve regularity by discretizing
@@ IV.3 Choice of time step
for some $C_{2}\lesssim 1$. In all our gauges $R_{,u}=0$ at $R=0$, and so (77) and (78) with $C_{1}=C_{2}$ are equivalent at the centre.
Since Paper II, we have realised that using $\Delta u=\Delta u_{1}$ may be an unnecessarily small timestep. We have found that $\Delta u=\Delta u_{2}$ with $C_{2}=0.1$ also leads to stable evolutions up to close to the apparent horizon, but with fewer timesteps overall.
@@
If we want to evolve through a horizon, where $R_{,x}=0$, we cannot use the numerical time step $\Delta u_{1}$. In double-null gauge $B=0$, the time step $\Delta u_{2}$ is not defined. The time steps $\Delta u_{3}$ and $\Delta u_{4}$ cannot be used for convergence testing as $\Delta u_{3}$ is independent of $\Delta x$. As an alternative to all these, for example if we want to test convergence of a solution in double-null gauge where $R_{,x}=0$ occurs, we also use the simple fixed timestep